# Adversarial Flow Models

### v1

@@ 4.2 Ablation Studies on the Hyperparameters
Table 1 shows a grid search of the optimal initial $\lambda_{ot}$ and $\lambda_{gp}$. Without OT loss, the training diverges regardless of $\lambda_{gp}$, proving the importance of the OT objective on stabilizing adversarial training for transformers. We also observe the phenomenon that an overly small $\lambda_{ot}$ fails to regularize the model and an overly large $\lambda_{ot}$ hurts distribution matching. Table 2 shows that decaying $\lambda_{ot}$ over time is critical for reaching the best FID.
@@
Table 3 shows that flow-based CG has a minor improvement over CG only at $t^{\prime}=0$. The optimal range of $\mathcal{U}(0,0.1)$ is much smaller than typical flow matching, likely because adversarial models already produce good samples without guidance.
@@ 4.3 Comparisons to the State of the Art
Footnote 5 compares single-step generation with guidance against state-of-the-art consistency-based and GAN models. Under the exact latent space and architectures, denoted by (•), our method achieves new best FIDs with large improvement margins across all model sizes, even when compared to concurrent works [101, 23, 91]. Notably, our B/2 model surpasses many XL/2 consistency-based models, likely because B/2 models are tight in capacity, and our method does not waste capacity on other timesteps. Our XL/2 model reaches a new best FID of 2.38. Table 8 shows the effect of guidance, and our method with only CG is still the best. Note that StyleGAN-XL has a slightly better FID while being smaller, but it operates in the pixel space, while ours is restricted by the VAE and DiT patch size 2. We mainly compare our method against others with the same settings.
@@
Footnote 5 shows that our model also achieves better FIDs in the few-step setting. Ours are trained on designated timesteps to conserve capacity. We find that any-step training performs worse due to the dilution of capacity and batch size. This is also observed in our 1D experiments (Fig. 1) where adversarial flow models need a larger batch size and converge more slowly for any-step training. This is often not a real limitation in practice, as achieving the best performance in a designated few-step setting is the priority.
@@
Footnote 5 shows that without guidance, even our 1NFE and 2NFE models outperform flow matching using 250NFE+ in the same latent space setting denoted by (•). We leave the exploration of combining adversarial flow models with representational methods to future work. However, we are concerned that current representational methods may have an unfair advantage. Since ImageNet is a small dataset where each class only contains about 1000 images, models cannot learn the correct generalizations from limited data on classes with large variations. However, DINO [5, 60] is trained on a much larger dataset, and using it may leak prior knowledge.
Extra-deep models.
Table 7 shows that our 56-layer and 112-layer models achieve better FIDs of 2.08 and 1.94, surpassing the 2 and 4-step counterparts. This confirms our hypothesis in Sec. 3.8. The results reveal an important insight: the quality of single-step generation may not be bounded by the training method, but by the depth of the generator. Depth scaling is a worthy direction for future research.
Computation efficiency.
[…]
@@
• iCT-XL/2 [79] 675M - None 1 34.24
• Shortcut-XL/2 [12] 675M 250 CFG [19] 1 10.60
• MeanFlow-B/2 [14] 131M 240 CFG [19] 1 6.17
• AlphaFlow-B/2 [101]  131M 240 CFG [19] 1 5.40
• MeanFlow-M/2 [14] 308M 240 CFG [19] 1 5.01
• MeanFlow-L/2 [14] 459M 240 CFG [19] 1 3.84
• MeanFlow-XL/2 [14] 676M 240 CFG [19] 1 3.43
TiM-XL/2 [91]  664M 300 CFG [19] 1 3.26
• AlphaFlow-XL/2 [101]  676M 240 CFG [19] 1 2.81
@@
BigGAN [4] 112M - cGAN [58] 1 6.95
GigaGAN [26] 569M 480 Match-loss [26] 1 3.45
GAT-XL/2+REPA [23]  602M 40 DA + cGAN 1 2.96
StyleGAN-XL [73] 166M - CG + cGAN 1 2.30
@@
• AF-B/2 130M 200 CG[9]+DA[30] 1 3.05
• AF-M/2 306M 120 CG[9]+DA[30] 1 2.82
• AF-L/2 457M 120 CG[9]+DA[30] 1 2.63
• AF-XL/2 673M 125 CG[9]+DA[30] 1 2.38
Table 4: Single-step generation on ImageNet 256px. [^1] In tables, (•) use the same latent space and architecture with parameters only differing by the number of timestep embeddings. [^2] In tables, ( ) use the same latent space but different architecture. [^4] In tables, ($\dagger$) are concurrent methods. [^5] In tables, iCT [79] results are reported by iMM [107].
@@
• IMM-XL/2 [107] 675M - CFG $\times2$ 3.99
TiM-XL/2 [91]  664M 300 CFG $\times2$ 3.61
• MeanFlow-XL/2 [14] 676M 240 CFG 2 2.93
• MeanFlow-XL/2 [14] 676M 1000 CFG 2 2.20
• AlphaFlow-XL/2 [101]  676M 240 CFG 2 2.16
@@
• AF-XL/2 675M 95 CG + DA 2 2.11
• AF-XL/2 675M 145 CG + DA 4 2.02
Table 5: Few-step generation on ImageNet 256px. [^1] In tables, (•) use the same latent space and architecture with parameters only differing by the number of timestep embeddings. [^2] In tables, ( ) use the same latent space but different architecture. [^4] In tables, ($\dagger$) are concurrent methods. [^5] In tables, iCT [79] results are reported by iMM [107].
@@
ADM [9] 554M 400 None 250 10.94
• DiT-XL/2 [62] 675M 1400 None 250 9.62
• SiT-XL/2 [56] 675M 1400 None 250 8.30
• SiT-XL/2+Disperse [89] 675M 1200 None 500 7.43
• SiT-XL/2+REPA [99] 675M 800 None 250 5.90
$\mathcal{R}$ RAE-XL [103] 676M 800 None 250 1.87
$\mathcal{R}$ SiT-XL/2+REPA-E [41] 675M 800 None 250 1.69
@@
MaskGIT [79] 227M 300 None 8 6.18
VAR [84] 310M 350 None 10 4.95
@@
• AF-B/2 130M 170 None 1 6.07
• AF-M/2 306M 110 None 1 5.21
• AF-L/2 457M 110 None 1 4.36
• AF-XL/2 673M 120 None 1 3.98
• AF-XL/2 675M 90 None 2 2.36
Table 6: No-guidance generation on ImageNet 256px. [^1] In tables, (•) use the same latent space and architecture with parameters only differing by the number of timestep embeddings. [^2] In tables, ( ) use the same latent space but different architecture. [^3] In tables, ( $\mathcal{R}$ ) use representational latent space. [^5] In tables, iCT [79] results are reported by iMM [107].
@@
Table 7: Deep architectures on ImageNet 256px.
@@
AF-XL/2 673M 120 None 1 3.98
AF-XL/2 673M 125 DA 1 3.86
AF-XL/2 673M 125 CG 1 2.54
AF-XL/2 673M 125 CG + DA 1 2.38
Table 8: Guidance type comparison on ImageNet 256px.
Additional content.
In the appendix, we discuss that our use of CG does not cause data leakage [39]. We also share more details, results, and visualizations.
Limitations.
Techniques in Section 3.9 can be improved with more principled methods. Current adversarial training uses more compute per iteration. $D$ network increases memory consumption. We use CG instead of CFG. We aim to address these in future work.
@@ 5 Conclusion
Our work proposes a framework to unify adversarial and flow modeling. We show that learning a deterministic transport greatly stabilizes training. We propose techniques to normalize gradients and incorporate guidance. Our method achieves new best FIDs and demonstrates training on deep architectures. Despite limitations, this offers exciting prospects for future research.
@@ Acknowledgement
We thank Chaorui Deng, Peng Wang, and Qing Yan for their valuable discussions on the methodology and manuscript preparation. We also thank Yang Zhao and Qi Zhao for their support in developing the dataloading and evaluation infrastructure.
Appendix A Qualitative Results
Qualitative results are provided below. Samples are generated using the same seeds across models.
@@ Appendix A Qualitative Results
Deterministic transport behavior is visible, particularly on bald eagle (class 22, first row) and geyser (class 974, last row), where the sky background (blue or white) is usually consistent on the same seed across models. However, details still vary across models. This is expected because (1) different model sizes have different degrees of generalization, (2) minibatch training has stochasticity, and (3) the OT scale is reduced toward the end of the training.
@@
Figure 9 shows comparisons with SiT [56]. We show that the adversarial objective generates perceptually better-looking samples even without guidance, while the flow matching method generates more perceptually out-of-distribution samples without guidance.
@@
Figure 9: Uncurated comparisons with SiT [56]. Same seeds. SiT uses Euler sampler. Classes are (207) golden retriever, (360) otter, (387) red panda, (974) geyser, (88) macaw, (979) valley, (417) balloon, (279) white fox, (22) bald eagle, (27) eft, (29) axolotl, (63) Indian cobra, (89) cockatoo, (360) otter, (972) cliff, (992) agaric.
@@
We follow prior works to project every layer’s hidden features onto the top-3 PCA components [23], and through a trained linear projection and decode to images using the VAE [47]. Note that in PCA, because DiT [62] uses absolute positional encoding through input, some PCA of early layers are dominated by sinusoidal encoding. Also, because the PCA of each layer is fitted independently, some layers have a different top-3 order, so the color of the visualization can change abruptly despite the underlying features being similar. Clear imagery is formed in the later layers of the model. Unlike [23], we do not impose any manual supervision losses on the intermediate features, and the models still obtain top FIDs regardless. Notice that for the 112-layer model, a large number of middle layers seem not to be contributing much in the visualization, but they are indeed effective in improving the final FID. The visualizations may not reveal the full contributions of these layers.
@@
The top row is hidden features at every layer projected onto the top-3 PCA components [23].
The bottom row is through a trained linear projection layer and decoded by the VAE [47].
@@
The top row is hidden features at every layer projected onto the top-3 PCA components [23].
The bottom row is through a trained linear projection layer and decoded by the VAE [47].
Appendix B Experimental Details

### v2

@@ 4.2 Ablation Studies on the Hyperparameters
Table 1 shows a grid search of the optimal initial $\lambda_{ot}$ and $\lambda_{gp}$. Without the OT loss, training diverges regardless of $\lambda_{gp}$, demonstrating the importance of the OT objective for stabilizing adversarial training in transformers. We also observe that an overly small $\lambda_{ot}$ fails to regularize the model, whereas an overly large $\lambda_{ot}$ hurts distribution matching. Table 2 shows that decaying $\lambda_{ot}$ over time is critical for reaching the best FID. In Table 11 of Appendix D, we show that the terminal OT scale can be further lowered given reduced learning rate.
@@
Table 3 shows that flow-based CG offers a modest improvement over CG applied only at $t^{\prime}=0$. The optimal range, $\mathcal{U}(0,0.1)$, is much smaller than in typical flow matching, likely because adversarial models already produce good samples without guidance.
@@ 4.3 Comparisons to the State of the Art
Footnote 5 compares single-step generation with guidance against state-of-the-art consistency-based and GAN models. Under the exact latent space and architectures, denoted by (•), our method achieves new best FIDs with large margins across all model sizes, even when compared to concurrent works (Zhang et al., 2025; Hyun et al., 2025; Wang et al., 2025). Notably, our B/2 model surpasses many XL/2 consistency-based models, likely because B/2 models are capacity-limited and our method does not waste capacity on other timesteps. Our XL/2 model reaches a new best FID of 2.38. Table 8 shows the effect of guidance, and our method with only classifier guidance (CG) and without discriminator augmentation (DA) is still the best. Note that StyleGAN-XL has a slightly better FID while being smaller, but it operates in the pixel space, while ours is restricted by the VAE and DiT patch size 2. We mainly compare our method against others with the same settings denoted by (•).
@@
Footnote 5 shows that our model also achieves better FIDs in the few-step setting. Our models are trained on designated timesteps to preserve capacity. We find that any-step training performs worse due to the dilution of capacity and batch size. This is also observed in our 1D experiments (Figure 1) where adversarial flow models need a larger batch size and converge more slowly for any-step training. In practice, this is often not a limitation, as achieving the best performance in a designated few-step setting is the priority.
@@
Footnote 5 shows that, without guidance, even our 1NFE and 2NFE models outperform flow matching with 250NFE+ in the same latent-space setting denoted by (•). This is due to the properties explained in Section 3.6.
Table 4: Single-step generation on ImageNet 256px. [^1] In tables, (•) use the same latent space and architecture with parameters only differing by the number of timestep embeddings. [^2] In tables, ( ) use the same latent but different architecture. [^4] In tables, ($\dagger$) are concurrent methods. [^5] In tables, iCT (Song and Dhariwal, 2023) results are reported by iMM (Zhou et al., 2025).
@@
• iCT-XL/2 675M - None 1 34.24
• Shortcut-XL/2 675M 250 CFG 1 10.60
• MeanFlow-B/2 131M 240 CFG 1 6.17
• AlphaFlow-B/2  131M 240 CFG 1 5.40
• MeanFlow-M/2 308M 240 CFG 1 5.01
• MeanFlow-L/2 459M 240 CFG 1 3.84
• MeanFlow-XL/2 676M 240 CFG 1 3.43
TiM-XL/2  664M 300 CFG 1 3.26
• AlphaFlow-XL/2  676M 240 CFG 1 2.81
@@
BigGAN 112M - cGAN 1 6.95
GigaGAN 569M 480 Match-loss 1 3.45
GAT-XL/2+REPA  602M 40 DA + cGAN 1 2.96
StyleGAN-XL 166M - CG + cGAN 1 2.30
@@
• AFM-B/2 130M 200 CG + DA 1 3.05
• AFM-M/2 306M 120 CG + DA 1 2.82
• AFM-L/2 457M 120 CG + DA 1 2.63
• AFM-XL/2 673M 125 CG + DA 1 2.38
Table 5: Few-step generation on ImageNet 256px. [^1] In tables, (•) use the same latent space and architecture with parameters only differing by the number of timestep embeddings. [^2] In tables, ( ) use the same latent but different architecture. [^4] In tables, ($\dagger$) are concurrent methods. [^5] In tables, iCT (Song and Dhariwal, 2023) results are reported by iMM (Zhou et al., 2025).
@@
• IMM-XL/2 675M - CFG $\times2$ 3.99
TiM-XL/2  664M 300 CFG $\times2$ 3.61
• MeanFlow-XL/2 676M 240 CFG 2 2.93
• MeanFlow-XL/2 676M 1000 CFG 2 2.20
• AlphaFlow-XL/2  676M 240 CFG 2 2.16
@@
• AFM-XL/2 675M 95 CG + DA 2 2.11
• AFM-XL/2 675M 145 CG + DA 4 2.02
Table 6: No-guidance generation on ImageNet 256px. [^1] In tables, (•) use the same latent space and architecture with parameters only differing by the number of timestep embeddings. [^2] In tables, ( ) use the same latent but different architecture. [^3] In tables, ( $\mathcal{R}$ ) use representational latent space. [^5] In tables, iCT (Song and Dhariwal, 2023) results are reported by iMM (Zhou et al., 2025).
@@
ADM 554M 400 None 250 10.94
• DiT-XL/2 675M 1400 None 250 9.62
• SiT-XL/2 675M 1400 None 250 8.30
• SiT-XL/2+Disperse 675M 1200 None 500 7.43
• SiT-XL/2+REPA 675M 800 None 250 5.90
$\mathcal{R}$ RAE-XL 676M 800 None 250 1.87
$\mathcal{R}$ SiT-XL/2+REPA-E 675M 800 None 250 1.69
@@
MaskGIT 227M 300 None 8 6.18
VAR 310M 350 None 10 4.95
@@
• AFM-B/2 130M 170 None 1 6.07
• AFM-M/2 306M 110 None 1 5.21
• AFM-L/2 457M 110 None 1 4.36
• AFM-XL/2 673M 120 None 1 3.98
• AFM-XL/2 675M 90 None 2 2.36
Table 7: Deep architectures on ImageNet 256px.
@@
Table 8: Guidance type comparison on ImageNet 256px.
@@
AFM-XL/2 673M 120 None 1 3.98
AFM-XL/2 673M 125 DA 1 3.86
AFM-XL/2 673M 125 CG 1 2.54
AFM-XL/2 673M 125 CG + DA 1 2.38
Extra-deep models.
Table 7 shows that our 56-layer and 112-layer models achieve improved FIDs of 2.08 and 1.94, surpassing their 28-layer 2-step and 4-step counterparts. This confirms our hypothesis in Section 3.8. The results reveal an important insight: the quality of single-step generation may not be bounded by the training method, but by the depth of the generator. Depth scaling is therefore a promising direction for future research.
4.4 Limitations and Future Works
Computation efficiency.
[…]
Additional limitations.
(1) $D$ network increases memory consumption. (2) We use CG instead of CFG. (3) Adversarial flow still requires techniques to mitigate the gradient vanishing problem (Appendix C).
Future works.
(1) The current adversarial flow models are discrete-time flow models. Extending it further to continuous-time flow modeling is a future direction. (2) Our work only explores training from scratch. We leave the exploration of post-training to future work.
@@ 5 Conclusion
Our work proposes a framework to combine adversarial and flow modeling. We show that learning a deterministic transport greatly stabilizes training. We propose techniques to normalize gradients and incorporate guidance. Our method achieves new best FIDs and demonstrates end-to-end training on 112-layer deep architectures. The framework and findings offer exciting prospects for future research.
@@ Acknowledgement
We thank Rohan Choudhury, Chaorui Deng, Peng Wang, and Qing Yan for their valuable discussions on the methodology and manuscript preparation. We also thank Yang Zhao and Qi Zhao for their support in developing the dataloading and evaluation infrastructure.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Quantitative Results
Table 9: Full metrics on conditional ImageNet 256px generation compared with other works. The table only shows methods and their configurations that are the most comparable to ours for comparison. (•) indicates methods using the same latent space and architecture as ours. Please refer to (Dhariwal and Nichol, 2021) for metric details.
[…]
Flow-matching and diffusion
ADM Pixel 554M 400 None 250 10.94 6.02 100.98 0.69 0.63
• DiT-XL/2 (Peebles and Xie, 2023) LDM (Rombach et al., 2022) 675M 1400 None 250 9.62 6.85 121.50 0.67 0.67
• SiT-XL/2 (Ma et al., 2024) LDM (Rombach et al., 2022) 675M 1400 None 250 8.30 - - - -
• SiT-XL/2+Disperse (Wang and He, 2025) LDM (Rombach et al., 2022) 675M 1200 None 500 7.43 - - - -
• SiT-XL/2+REPA (Yu et al., 2024) LDM (Rombach et al., 2022) 675M 800 None 250 5.90 - - - -
RAE-XL (Zheng et al., 2025a) DINOv2 (Oquab et al., 2023) 676M 800 None 250 1.87 - 209.70 0.80 0.63
SiT-XL/2+REPA-E (Leng et al., 2025) Joint-trained 675M 800 None 250 1.69 4.17 219.30 0.77 0.67
[…]
Consistency-based models
• Shortcut-XL/2 (Frans et al., 2024) LDM (Rombach et al., 2022) 675M 250 CFG 1 10.60 - - - -
TiM-XL/2 (Wang et al., 2025) 664M 300 None 1 7.11 4.97 140.39 0.71 0.63
• MeanFlow-B/2 (Geng et al., 2025) 131M 240 CFG 1 6.17 - - - -
• AlphaFlow-B/2 (Zhang et al., 2025) 131M 240 CFG 1 5.40 - - - -
• MeanFlow-M/2 (Geng et al., 2025) 308M 240 CFG 1 5.01 - - - -
• MeanFlow-L/2 (Geng et al., 2025) 459M 240 CFG 1 3.84 - - - -
• MeanFlow-XL/2 (Geng et al., 2025) 676M 240 CFG 1 3.43 - - - -
TiM-XL/2 (Wang et al., 2025) 664M 300 CFG 1 3.26 4.37 210.33 0.75 0.59
• AlphaFlow-XL/2 (Zhang et al., 2025) 676M 240 CFG 1 2.81 - - - -
• Shortcut-XL/2 (Frans et al., 2024) LDM (Rombach et al., 2022) 675M 250 CFG 4 7.80 - - - -
[…]
• IMM-XL/2 (Zhou et al., 2025) 675M - CFG $\times2$ 3.99 - - - -
TiM-XL/2 (Wang et al., 2025) 664M 300 CFG $\times2$ 3.61 6.74 151.79 0.74 0.59
• MeanFlow-XL/2 (Geng et al., 2025) 676M 1000 CFG 2 2.20 - - - -
• AlphaFlow-XL/2 (Zhang et al., 2025) 676M 240 CFG 2 2.16 - - - -
Autoregressive, masking, and hybrids
MaskGIT (Song and Dhariwal, 2023) VQGAN (Esser et al., 2021) 227M 300 None 8 6.18 - 182.10 0.80 0.51
VAR (Tian et al., 2024) MS-VQVAE (Tian et al., 2024) 310M 350 None 10 4.95 - - - -
[…]
GANs
BigGAN (Brock et al., 2018) Pixel 112M - cGAN 1 6.95 7.36 171.40 0.87 0.28
GigaGAN (Kang et al., 2023) Pixel 569M 480 Match-loss 1 3.45 - 225.52 0.84 0.61
GAT-XL/2+REPA (Hyun et al., 2025) LDM (Rombach et al., 2022) 602M 40 DA+cGAN 1 2.96 - - - -
StyleGAN-XL (Sauer et al., 2022) Pixel 166M - CG+cGAN 1 2.30 4.02 265.12 0.78 0.53
Adversarial flow models (Ours)
• AFM-B/2 LDM (Rombach et al., 2022) 130M 170 None 1 6.07 5.31 169.51 0.72 0.49
• AFM-M/2 306M 110 None 1 5.21 5.60 178.48 0.75 0.54
• AFM-L/2 457M 110 None 1 4.36 5.39 186.21 0.77 0.53
• AFM-XL/2 673M 120 None 1 3.98 5.40 201.85 0.78 0.52
• AFM-XL/2 673M 90 None 2 2.36 4.35 235.77 0.81 0.52
• AFM-B/2 LDM (Rombach et al., 2022) 130M 200 CG+DA 1 3.05 5.32 269.18 0.81 0.51
• AFM-M/2 306M 120 CG+DA 1 2.82 5.20 279.12 0.81 0.50
• AFM-L/2 457M 120 CG+DA 1 2.63 5.10 277.96 0.81 0.52
• AFM-XL/2 673M 125 CG+DA 1 2.38 4.87 284.18 0.81 0.52
• AFM-XL/2 675M 95 CG+DA 2 2.11 4.33 273.84 0.82 0.55
[…]
• AFM-XL/2 675M 145 CG+DA 4 2.03 4.59 259.66 0.78 0.59
[…]
AG (Karras et al., 2024), CFG (Ho and Salimans, 2022), CG (Dhariwal and Nichol, 2021), cGAN (Miyato and Koyama, 2018), DA (Karras et al., 2020a), Match-loss (Kang et al., 2023)
Appendix B Qualitative Results
Qualitative results are provided below. Samples are generated with the same seeds across models.
@@ Appendix B Qualitative Results
Deterministic transport behavior is visible, particularly on bald eagle (class 22, first row) and geyser (class 974, last row), where the sky background (blue or white) is usually consistent on the same seed across models. However, details still vary from model to model. This is expected because (1) different model sizes have different degrees of generalization, (2) minibatch training has stochasticity, and (3) the OT scale is reduced toward the end of training.
@@
Figure 9 shows comparisons against SiT (Ma et al., 2024). We show that the adversarial objective generates perceptually better-looking samples even without guidance, while the flow-matching method generates more perceptually out-of-distribution samples without guidance.
@@
Figure 9: Uncurated comparisons with SiT (Ma et al., 2024). Same seeds. SiT uses Euler sampler. Classes are (207) golden retriever, (360) otter, (387) red panda, (974) geyser, (88) macaw, (979) valley, (417) balloon, (279) white fox, (22) bald eagle, (27) eft, (29) axolotl, (63) Indian cobra, (89) cockatoo, (360) otter, (972) cliff, (992) agaric.
@@
We follow prior works to project the hidden features of every layer onto the top three PCA components (Hyun et al., 2025), and through a trained linear projection and decode them into images using the VAE (Lin et al., 2025a). Note that in PCA, because DiT (Peebles and Xie, 2023) uses absolute positional encoding through input, the PCA of some early layers is dominated by sinusoidal encoding. Also, because the PCA of each layer is fitted independently, some layers have a different top-three ordering, so the color of the visualization can change abruptly despite the underlying features being similar. Clear imagery is formed in the later layers of the model. Unlike (Hyun et al., 2025), we do not impose any manual supervision losses on the intermediate features, and the models still obtain top FIDs. Notice that for the 112-layer model, a large number of middle layers seem not to be contributing much in the visualization, but they are indeed effective in improving the final FID. The visualizations may not reveal the full contributions of these layers.
@@
The top row is hidden features at every layer projected onto the top-3 PCA components (Hyun et al., 2025).
The bottom row is through a trained linear projection layer and decoded by the VAE (Lin et al., 2025a).
@@
The top row is hidden features at every layer projected onto the top-3 PCA components (Hyun et al., 2025).
The bottom row is through a trained linear projection layer and decoded by the VAE (Lin et al., 2025a).
Appendix C Techniques for Adversarial Training
We share additional training techniques that, while imperfect, are very effective. They address the remaining challenges of adversarial training, not introduced by our adversarial flow models.
On optimization, minimax optimization with gradient descent is prone to oscillation. The equilibrium is better achieved by the weight average than at the last iterate (Daskalakis et al., 2017). We find the optimistic optimizer (Daskalakis et al., 2017) and asymmetrical learning rates (Heusel et al., 2017) ineffective. Our approach is to keep an EMA of $G$ (Yazıcı et al., 2018), which consistently outperforms the online model by a large margin. More importantly, once the performance of the EMA model plateaus toward the end of training, we replace online $G$ with the EMA weights and repeat this procedure automatically after every subsequent epoch. The learning rate is also decreased during this phase. We find this simple technique very effective for approaching peak performance.
On training dynamics, many techniques have been proposed to address the vanishing gradient problem. WGAN (Arjovsky et al., 2017) proposes learning the Wasserstein-1 distance but requires a K-Lipschitz $D$. DiffusionGAN (Wang et al., 2022) projects $D$ onto a flow process in the same spirit as our approach with guidance. Discriminator augmentation (DA) (Karras et al., 2020a) is another approach to increase the support overlap. It has been used by recent GANs (Huang et al., 2024; Hyun et al., 2025). However, we believe that the choice of augmentation may implicitly inject inductive biases. For example, prior work finds that affine transforms outperform other distortions on image data (Karras et al., 2020a) because it implicitly encourages $D$ to recognize affine transforms as a more acceptable generalization for image generation. The last approach is to simply reload $D$ from an earlier checkpoint to reset the pace. Our experiments find that reloading $D$ performs surprisingly well, whereas DiffusionGAN is less effective in our preliminary studies. Therefore, we reload $D$ when training stalls for the no-guidance setting to avoid introducing any additional biases, and additionally use DA for the guidance setting.
Note that our work mainly focuses on stabilizing the generator through learning a deterministic transport map, generalizing adversarial training to flow modeling, and introducing the support for guidance. Our work does not address the gradient vanishing problem and relies on techniques proposed by previous research.
Appendix D Experimental Details